OpenStudy (anonymous):

hi! could someone tell me how to convert kg to g, or cm to m. I can figuire out the answer just not sure how to actually solve it. Is there a different equation for converting all metric units? or one for all?

OpenStudy (australopithecus):

1kg = 1000g 1 m = 100cm

OpenStudy (australopithecus):

if you want to convert 2.5grams to kg (2.5g)(1kg/1000g) notice how g cancels out and you are left with kg

OpenStudy (australopithecus):

same thing for cm 55m(100cm/1m) meters cancel out and you are left with cm

OpenStudy (australopithecus):

(55*100)m*cm/m = (55*100)cm

OpenStudy (australopithecus):

hope that makes sense?

OpenStudy (anonymous):

okay so if i have 3.68kg and i want to convert it to g. I would multiply 1000 times 3.68? to get 3,680. then would i divide it?

OpenStudy (australopithecus):

when you are dealing with most units this is how it works temperature is a bit different

OpenStudy (australopithecus):

\[\frac{3.68kg}{1}*\frac{1000g}{1kg} = \frac{3680g*kg}{1*1kg} = \frac{3680g*kg}{kg} = \frac{3680g*(1)}{1} = \frac{3680g}{1} = 3680g\]

OpenStudy (australopithecus):

remember you can divide anything by 1 it doesn't change its value, in fact every real number can be represented with 1 as its denominator

OpenStudy (australopithecus):

hope that fact helps you with fractions, just treat the units kind of like numbers (even though they aren't)

OpenStudy (anonymous):

thankyou! :)